please help me to solve this..........
tanA+tanB+tanC=0 , A+B+C=180 , cotA+cotB+cotC=?

Grade:11

1 Answers

ABHISHAY TURNER
6 years ago
Given- A+B+C=180
=>A+B=180 – C
MULTIPLYING COT ON BOTH SIDES
=>COT(A+B)=COT(180-C)                                                          
=>COT(A+B)=                             ––––––( 1 )
 
=>COT(A+B)= 1
     – COT C   
=> – COT C = 1
     – COT C  
HENCE,
COT A+ COT B+ COT C=1